# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v3

@@ 4.1 Hierarchical Data Representation
Figure 3: HiSL module for intra-fact and inter-fact message passing.
4.2 Hierarchical Structure Learning
@@
It’s evident that HiDR form introduces many additional relation nodes and fact nodes. To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Fig. 3.
@@
We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{HiDR}|\times d}$ for atomic nodes and all relation edges. Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$ with a projection matrix $\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$: $\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}}$, where $\mathbf{E_{a}}\subseteq\mathbf{E}$ is the atomic relation embeddings. Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
@@
where $(h,r,t)\in\mathcal{F}_{a}$, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is the concatenation operation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding and $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP. Therefore, the initialization of relation nodes and fact nodes is sufficiently parameter-efficient.
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec [18]:
@@
where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as periodic function, $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP as non-periodic function, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
@@
In this stage, massage passing is conducted for fact nodes. Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$. Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{HiDR}$. For this subgraph, we employ the graph attention network [5] to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$. The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
@@
$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$ (4)
@@
At this stage, message passing is conducted on the whole graph $\mathcal{G}^{HiDR}$. Similar to previous work [31], we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges. We employ the circular-correlation operator, defined as:
$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$ (5)
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, the $\odot$ is the element-wise (Hadamard) product. Furthermore, in order to fully capture the heterogeneity of the graph, we classify edges along two dimensions: $\lambda(r)\in\left\{forward,reverse\right\}$ and $\tau(r)\in\{connected\,\,relation,atomic\,\,$
$relation,nested\,relation\}$ and adopt two relation-type specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for more fine-grained aggregation:
$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$ (6)
$\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l},$ (7)
where $\mathbf{W}^{l}_{self},\mathbf{W}^{l}_{rel}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid activation function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$. We utilize $\phi\left(\cdot\right)$ to combine the information from edge $r$ and node $j$, and then passes it to node $i$ for update.
Through Intra-fact and Inter-fact two-stage message passing, nodes can fully capture both local semantic and global structural information. Moreover, its number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
4.3 Link Prediction Decoder
Since the query varies across different settings, we use the transformer [32] as the decoder with the mask pattern. Specifically, we convert the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input to the transformer. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
$P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right),$ (8)
@@ 4.3 Link Prediction Decoder
where $P\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $[\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes. The $P_{t}$ and $y_{t}$ are probability and ground truth of the $t$-th candidate. The final loss function $\mathcal{L}$ includes both node loss and edge loss during the predictions.
5 Experiment
5.1 Experiment Settings

### v4

@@ Introduction
Figure 3: The overview of UniHR, including HiSL module for intra-fact and inter-fact message passing.
Next, we introduce how to transform different types of beyond-triple KGs into HiDR form.
For HKGs, we regard the key-value pairs as complementary information for facts. Thus, we translate the H-Facts $\mathcal{F}_{\text{HKG}}\,\text{=}\{((h,r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m}\}$ into the HiDR form that $\mathcal{G}^{\text{HiDR}}_{\text{HKG}}\,\text{=}\,\{\mathcal{V},\mathcal{R},\mathcal{F}^{\text{HiDR}}_{\text{HKG}}\}$ following the definition, where $\mathcal{F}_{c}\,\text{=}$
$\{(f,has\,relation,e_{r}),(f,has\,head\,entity,h),(f,has\,tail\,$
$entity,t),(f,k_{1},v_{1}),\ldots,(f,k_{m},v_{m})\}$, $\mathcal{F}_{a}\text{=}\{(h,r,t)\,|((h,$
$r,t),\{(k_{i}\text{:}\,v_{i})\}_{i=1}^{m})\in\mathcal{F}_{\text{HKG}})\}$ and $\mathcal{F}_{n}\text{=}\,\varnothing$.
For NKGs, HiDR can naturally represent the hierarchical facts, so we translate the atomic facts $\mathcal{F}_{\text{NKG}}\,\text{=}\,\{(h_{i},r_{i},t_{i})\}$ and the nested facts $\hat{\mathcal{F}}_{\text{NKG}}\,\text{=}\,\{((h_{1},r_{1},t_{1}),R,(h_{2},r_{2},t_{2}))|$
$(h_{i},r_{i},t_{i})\in\mathcal{F}_{\text{NKG}}\}$ into the form of HiDR that $\mathcal{G}^{\text{HiDR}}_{\text{NKG}}\,\text{=}\,\{\mathcal{V},$
$\mathcal{R},\mathcal{F}^{\text{HiDR}}_{\text{NKG}}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h_{i},r_{i},t_{i})|$
$(h_{i},r_{i},t_{i})\in\mathcal{F}_{\text{NKG}}\}$, $\mathcal{F}_{c}\,\text{=}\{(f_{i},has\,head\,entity,h_{i}),(f_{i},$
$has\,tail\,entity,t_{i}),(f_{i},has\,relation,e_{r_{i}})|f_{i}\,\text{=}\,(h_{i},r_{i},t_{i})\in$ $\mathcal{F}_{\text{NKG}}\}$ and $\mathcal{F}_{n}\text{=}\{(f_{1},R,f_{2})|f_{i}\in\mathcal{F}_{\text{NKG}}\}$.
For TKGs, we regard the TKG as a special HKG, and convert timestamps to auxiliary key-value pairs in HKGs by adding two special atomic relations: begin and end, regarding timestamps as special numerical atomic nodes. Thus, we firstly translate the temporal facts in TKGs $\mathcal{F}_{\text{TKG}}\text{=}\,\{(h,r,t,[\tau_{b},\tau_{e}])\}$ into H-Facts form $\mathcal{F}_{\text{TKG}}^{\text{HKG}}\text{=}\,\{(h,r,$
$t,begin\text{:}\tau_{b},end\text{:}\tau_{e})\}$. Then according to the previous transformation in HKG, it can be translated into the HiDR form that $\mathcal{G}_{\text{TKG}}^{\text{HiDR}}\text{=}\{\mathcal{V},\mathcal{R},\mathcal{F}_{\text{TKG}}^{\text{HiDR}}\}$ following the definition, where $\mathcal{F}_{a}\,\text{=}\,\{(h,r,t)\,|\,(h,r,t,begin\text{:}\,\tau_{b},end\text{:}\tau_{e})\in\mathcal{F}_{\text{TKG}}^{\text{HKG}}\}$, $\mathcal{F}_{c}\,\text{=}$
$\{(f,has\,relation,e_{r}),(f,has\,head\,entity,h),\,(f,has\,tail\,$
$entity,t),\,(f,begin,\tau_{b}),(f,end,\tau_{e})\,|\,f\,\text{=}\,(h,r,t,begin\text{:}\,\tau_{b},$
$end\text{:}\,\tau_{e})\in\mathcal{F}_{\text{TKG}}^{\text{HKG}}\}$ and $\mathcal{F}_{n}\,\text{=}\,\varnothing$.
In summary, HiDR serves as a module that dynamically transforms the original data into a unified representation optimized for the model, without altering the storage format of the original data. Moreover, from the perspective of graph learning, it fully preserves the semantics of the original knowledge graphs without any loss of information.
Hierarchical Structure Learning
@@
It’s evident that HiDR form introduces many additional relation nodes and fact nodes. To avoid significantly increasing the model’s training parameters while fully capturing the hierarchy of HiDR form, we design a Hierarchical Structure Learning module, abbreviated as HiSL shown in Figure 3.
@@
We first initialize the embedding matrices $\mathbf{H}_{a}\in\mathbb{R}^{|\mathcal{V}_{a}|\times d}$ and $\mathbf{E}\in\mathbb{R}^{|\mathcal{R}^{\text{HiDR}}|\times d}$ for atomic nodes and all relation edges. Then we also initialize the embedding of relation node $\mathbf{H}_{r}\in\mathbb{R}^{|\mathcal{V}_{r}|\times d}$, which can be transformed from the relation edge $r$ with a projection matrix $\mathbf{W_{r}}\in\mathbb{R}^{d\times d}$: $\mathbf{H}_{r}=\mathbf{E_{a}\cdot W_{r}}$, where $\mathbf{E_{a}}\subseteq\mathbf{E}$ is the atomic relation embeddings. Then we initialize the fact node embeddings $\mathbf{H}_{f}$ to explicitly capture key information within facts by utilizing the embedding of (main) triple:
@@
where $(h,r,t)\in\mathcal{F}_{a}$, $f_{m}\text{:}\;\mathbb{R}^{3d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, $\cdot$$\cdot$] is the concatenation, $\mathbf{h}_{h},\mathbf{h}_{t}\subseteq\mathbf{H}_{a},\mathbf{h}_{r}\subseteq\mathbf{H}_{r}$ denote (main) triple embedding. Therefore, the initialization of relation nodes and fact nodes is sufficiently parameter-efficient.
For numerical atomic nodes, namely timestamps in temporal knowledge graphs, we encode the timestamp $\tau$ into an embedding with Time2Vec (Kazemi et al. 2019):
@@
where $f_{p}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ and $f_{np}\text{:}\;\mathbb{R}^{1}\rightarrow\mathbb{R}^{d}$ are both 1-layer MLPs as periodic and non-periodic functions, and $\omega_{p}\in\mathbb{R}^{1}$ is a learnable parameter for scaling the periodic features.
@@
In this stage, massage passing is conducted for fact nodes. Given a fact node $f_{k}\in\mathcal{V}_{f}$, we construct its constituent elements, i.e., one-hop neighbors, as the node set $\mathcal{V}_{k}\,\text{=}\,\{v\in\mathcal{N}_{f_{k}}\,|\,v\in\mathcal{V}_{a}\cup\mathcal{V}_{r}\}$, where $\mathcal{N}_{f_{k}}$ is the set of one-hop neighbors of fact node $f_{k}$. Then we retain the edges directly connected to fact node $f_{k}$, thereby constructing a subgraph $\mathcal{G}_{k}\,\text{=}\,\{\mathcal{V}_{k},\mathcal{R}_{k},\mathcal{F}_{k}\}\subseteq\mathcal{G}^{\text{HiDR}}$.
WikiPeople WD50K
Model subject/object all entities subject/object all entities
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
NaLP (Guan et al. 2019) 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
StarE (Galkin et al. 2020) 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
GRAN (Wang et al. 2021) 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
tNaLP (Guan et al. 2023) 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
HyNT (Chung, Lee, and Whang 2023) 0.479 0.411 0.601 0.478 0.409 0.601 0.328 0.256 0.468 0.356 0.285 0.493
ShrinkE (Xiong et al. 2023a) 0.485 0.431 0.601 - - - 0.345 0.275 0.482 - - -
HAHE (Luo et al. 2023) 0.498 0.418 0.610 0.497 0.421 0.614 0.343 0.269 0.484 0.378 0.306 0.515
NYLON (Yu, Yang, and Yang 2024) 0.385 0.299 0.527 0.384 0.300 0.520 0.326 0.262 0.446 0.291 0.226 0.414
HyperSAT (Wang, Chen, and Zhang 2025) 0.493 0.427 0.610 0.496 0.430 0.613 0.345 0.270 0.489 0.380 0.306 0.520
UniHR 0.496 0.419 0.619 0.496 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.517
Table 1: Results on HKG datasets, are reproduced by us and others are taken from (Wang, Chen, and Zhang 2025).
For this subgraph, we employ the graph attention to aggregate local information, computing the attention score $\alpha_{i,j}$ between node $i\in\mathcal{V}_{k}$ and its neighbor $j$. The formula for calculating $\alpha_{i,j}$ in the $l$-th layer is as follows:
@@
$\mathbf{h}_{i}^{l}=\mathbf{h}_{i}^{l}+\sum\limits_{j\in\mathcal{N}_{i}}\alpha_{i,j}^{l}\cdot\mathbf{W}_{out}^{l}\mathbf{h}_{j}^{l}.$
@@
At this stage, message passing is conducted on the whole graph $\mathcal{G}^{\text{HiDR}}$. Specifically, we use a non-parametric aggregation operator $\phi\left(\cdot\right)\text{:}\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to obtain messages from neighbouring nodes and edges. We employ the circular-correlation operator, defined as:
$\phi\left(\mathbf{h}_{j},\mathbf{e}_{r}\right)=\mathbf{h}_{j}\star\mathbf{e}_{r}=\mathbf{F}^{-1}\left(\left(\mathbf{F}\mathbf{h}_{j}\right)\odot\overline{\left(\mathbf{F}\mathbf{e}_{r}\right)}\right),$
where $\mathbf{F}$ and $\mathbf{F}^{-1}$ denote the discrete fourier transform (DFT) matrix and its inverse matrix, and the $\odot$ is the element-wise product. In order to fully capture the graph’s heterogeneity, we classify edges along two dimensions: direction $\lambda(r)\in\left\{forward,reverse\right\}$ and type $\tau(r)\in\{connected\,\,relation,atomic\,\,relation,nested\,\,relation\}$ and adopt two relation-specific learnable parameters $\mathbf{W}_{\lambda\left(r\right)}$
$\in\mathbb{R}^{d\times d}$ and $\omega_{\tau\left(r\right)}\in$ $\mathbb{R}^{1}$ for fine-grained aggregation:
$\mathbf{h}_{i}^{l+1}=\sum\limits_{(r,j)\in\mathcal{N}(i)}{\sigma\left(\omega_{\tau\left(r\right)}^{l}\right)\mathbf{W}_{{\lambda}\left(r\right)}^{{l}}\phi\left(\mathbf{h}_{j}^{l},\mathbf{e}_{r}^{l}\right)}+\mathbf{W}_{self}^{l}\mathbf{h}_{i}^{l},$
where $\mathbf{W}^{l}_{self}\in\mathbb{R}^{d\times d}$, $\sigma$ is a sigmoid function and $\mathcal{N}\left(i\right)$ is a set of immediate neighbors of $i$ for its outgoing edges $r$. We utilize $\phi\left(\cdot\right)$ to combine the information from edge $r$ and node $j$, and then passes it to node $i$ for update. Meanwhile, we update the relation representation as: $\mathbf{e}_{r}^{l+1}=\mathbf{W}_{rel}^{l}\mathbf{e}_{r}^{l}$.
Through Intra-fact and Inter-fact two-stage message passing, nodes can fully capture both local semantic and global structural information. Moreover, the number of training parameters does not increase with the scale of the graph, thereby effectively adapting to the HiDR form.
Link Prediction Decoder
Since the query varies across different settings, we use the transformer (Vaswani et al. 2017) as the decoder with mask pattern. Specifically, we serialize the updated node and edge embeddings into a sequence of fact embeddings, mask the elements to be predicted in facts with the $\bm{\left[M\right]}$ token as the input. Finally, we obtain the embedding of output $\bm{\left[M\right]}$ in the last layer to measure the plausibility of the fact, denoted as $\mathbf{h}_{pre}$, and calculate the probability distribution of candidates, followed by training it using the cross-entropy loss:
@@
where $P=\,{\rm Softmax}\left(f\left(\mathbf{h}_{pre}\right)[\mathbf{E};\mathbf{H}]^{\top}\right)\in\mathbb{R}^{\left|\mathcal{R}\right|+\left|\mathcal{V}\right|}$ represents the confidence scores of all candidates, $f\text{:}\;\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ is a 1-layer MLP, and $\mathbf{E};\mathbf{H}]\in\mathbb{R}^{(\left|\mathcal{R}\right|+\left|\mathcal{V}\right|)\times d}$ is the embedding matrix of all candidate edges or nodes. The $P_{t}$ and $y_{t}$ are probability and ground truth of the $t$-th candidate.
Experiment
Experiment Settings